Let s n =∑k=1 to n[n/(n 2 +kn+k 2 )] and T n =∑k=0 to (n-1) [n/(n 2 +nk+k 2 )] for n=1,2,3....... Then A S n ‹ pi/3√3 C T n B Sn> pi/3 √3 D T n>pi /3√3

Let sn=∑k=1 to n[n/(n2+kn+k2)] and Tn=∑k=0 to (n-1)  [n/(n2+nk+k2)] for n=1,2,3....... Then


A  Sn‹ pi/3√3                          C  Tn

B  Sn> pi/3√3                           D  Tn>pi/3√3

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Hi vanya

You can find Sn by summation of series using definite intigral as the limit of sum

sn=∑k=1 to n[n/(n2+kn+k2)]

sn=1/n∑k=1 to n [n2/(n2+kn+k2)]

sn=1/n∑k=1 to n [1/(1+k/n+(k/n)2)]

sn =∫1/(1+x+x2)dx   limit 0 to 1

   after solving

sn =pi/3√3


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation. All the best vanya
 
 

Regards,

Askiitians Experts

BAdiuddin